OpenStudy (anonymous):

Name a number between 3/8 and 3/7 and how do you do it?

OpenStudy (blockcolder):

When you say "number", do you mean a rational number? If so, then just take the average of 3/8 and 3/7.

OpenStudy (lgbasallote):

if you just mean number between...then think of a number greater than 3/8 but less than 3/7

OpenStudy (anonymous):

This is a fifth grade math problem. Please explain the process.

OpenStudy (lgbasallote):

hmmm okay..to make it easier..change both fractions to common terms by that i mean numbers with common denominators...do you know how?

OpenStudy (anonymous):

I know that 56 is a common denominator.

OpenStudy (lgbasallote):

right! but what is 3/8 in terms of a 56 denominator? to get that...divide 56 by your original denominator (which is 8)..that will give you 7..now multiply that to your numerator (which is 3)..that will give you 21 so...3/8 = 21/56 try it with the other fraction

OpenStudy (anonymous):

24/56. Is this the simplest way to figure this out?

OpenStudy (lgbasallote):

yeppp...now you have 21/56 and 24/56 think of a number between those 2 numbers.. Hint: the denominator is 56

OpenStudy (anonymous):

23/56 or 22/56. Is there simpler way?

OpenStudy (lgbasallote):

afraid not..this is the simplest i can think of

OpenStudy (anonymous):

Thank you!

OpenStudy (lgbasallote):

<tips hat>

OpenStudy (anonymous):

There are as you know infinite number of solutions for your problem here some of them \[\left\{\frac{45}{112},\frac{23 }{56},\frac{93}{224},\frac{ 117}{280},\frac{47}{112},\frac{165}{392},\frac{27}{64}\\ ,\frac{71}{168},\frac{237}{ 560},\frac{261}{616},\frac{ 95}{224},\frac{309}{728},\frac{333}{784},\frac{17}{40}\\ ,\frac{381}{896},\frac{405} {952},\frac{143}{336},\frac {453}{1064},\frac{477}{1120 }\right\} \] The above ones were generated as a weighted averages \[\frac{3}{7} \left(1-\frac{1}{i}\right)+ i \frac{3}{8 },\,\, i=2,3,4, \cdots 20 \]